OpenStudy (anonymous):

What is the molar mass of (NH4)2O? Explain how you calculated this value.

OpenStudy (anonymous):

@JFraser @ganeshie8

OpenStudy (anonymous):

I think it is 52 g/mol is that correct?

OpenStudy (australopithecus):

http://www.wolframalpha.com/input/?i=%28NH4%292O+molar+mass yes you are right

OpenStudy (jotopia34):

@wriddler. I think its actually 68g/mol if this is what you mean. \[(NH _{4})_{2}O _{2}\] The explanation is : You take the first part (NH4)2. This says that there are 2 Nitrogens, at 14g/mol. = 28 (4H's) x 2. at H= 1g/mol, so 4*2*1=8. Total that and get 28+8=36. Next do the O2. This says there are 2 Oxygens. Each oxygen is 16g/mol so, 2*16=32 Add that up 36+32 and get 68g/mol

OpenStudy (anonymous):

but it said (NH4)2O, wouldnt that be like (NH4)2 + O essentially?

OpenStudy (anonymous):

disregard what I last said I just realized something

OpenStudy (anonymous):

the 2 in 2O is subtext

OpenStudy (jotopia34):

Do you mean \[(NH_{4})O _{2}\]

OpenStudy (anonymous):

\[(NH _{4})_{2}O\]

OpenStudy (anonymous):

thats what the question says

OpenStudy (anonymous):

The answer that I saw somewhere else was: N =14x2 = 28 H = 8 O = 16 Add them up 28 + 16 + 8 = 52, molar mass is written grams over moles or "g/mol" so 52g/mol but it doesnt seem very correct

OpenStudy (anonymous):

@Australopithecus What do you think?

OpenStudy (jotopia34):

I'm pretty sure as you have stated, you have 2 (NH4)s. That means two of each element inside the bracket Therefore 2 Nitrogens 4 hydrogens times 2 2N=28g/mol 8H= 8g/mol Add those 2 and get 28+8=36 Next do the oxygen. 1 oxygen is 16g/mol Add the 36+16 and you do get 52g/mol

OpenStudy (anonymous):

okay thanks, mind helping me with one more?

OpenStudy (jfraser):

\((NH_4)_2O\) certainly is a compound, it's ammonium oxide

OpenStudy (australopithecus):

yeah I'm just dumb :\

OpenStudy (australopithecus):

I should have known it exists durrrr